# Integrating representative and non-representative survey data for efficient inference

### v2

@@ 1 Introduction
If we draw our attention to the similarities between causal inference in observational studies and non-representative sampling, as suggested by Mercer et al. (2017), we can better understand the main assumptions required to obtain unbiased estimates. The key assumption for post-survey adjustments is that of “conditional ignorability” (Mercer et al. 2017, Schuessler & Selb 2019). Much like the assumption of “no unmeasured confounders” that is required to obtain unbiased causal estimates in an observational study, the conditional ignorability assumption cannot be guaranteed. In other words, there is no sure way of choosing a sufficient set of adjustment covariates (based on available data alone), and ultimately there is no test or way to validate when the assumption of conditional ignorability is met.
@@
In Section 2, we outline the proposed method, state all assumptions, and provide an example where including biased surveys will reduce the uncertainty of estimates. In Section 3, we conduct a simulation study to show that using conservative assumptions, mis-specifying the model is not overly detrimental, as long as the complexity of the model for the bias is proportional to the number of time-points. In Section 4, inspired by the dataset presented by Bradley et al. (2021), we show that the proposed method performs well in estimating the vaccination rate by combining two “biased” surveys with a presumably “unbiased” survey. The results show a noticeable reduction in uncertainty (compared to only using the unbiased survey). We conclude in Section 5 by discussing potential uses and limitations for using large amounts of unrepresentative survey data to improve the precision of survey estimates.
2 Methods
@@
The hyper-geometric distribution describes the probability of drawing $Y_{kt}$ positive individuals amongst a sample of $n_{kt}$ who responded to the survey (without any individuals responding more than once), from a finite population of size $N$ that contains exactly $P_{t}$ positive individuals. The non-central hyper-geometric is a generalization of the hyper-geometric distribution whereby, in this case, the response may be biased so that either positive or negative individuals are more likely to respond. When $\phi_{kt}>1$, positive individuals are more likely to respond than negative individuals; when $\phi_{k}<1$, positive individuals are less likely to respond than negative individuals. When $\phi_{k}=1$, we have that the probability of responding is equal for both positive and negative individuals, and the non-central hyper-geometric distribution reduces to the standard hyper-geometric distribution. In this parameterization, the $\phi_{kt}$ parameter can be interpreted as an odds ratio: the odds of a positive individual being sampled vs. the odds of a non-positive individual being sampled. For a more detailed breakdown of the non-central hyper-geometric distribution, its use in population size estimation using surveys, and a discussion of Bayesian estimation of its parameters see Ballerini & Liseo (2022).
@@
To illustrate the proposed method, consider the goal of performing inference for $T=10$ time-points and $K=3$ surveys, with a population of $N=10,000$. Survey 1 is unbiased, whereas survey 2 and 3 are compromised by selection bias. Specifically, for these two surveys we generate data under the assumption of a linear $\phi_{kt}$ (8). All the parameters and data are generated jointly using (1-3), (5-6, 8), and the priors used in analysing the dataset are as such. The data are illustrated in Table 1.
@@
Analysis of each survey in isolation assumes $\phi_{kt}=1$ for all $t$, but retain all the same priors (1-3), and (5-6). In Figure 1, we see that using the biased surveys (k = 2 & 3, assuming they are unbiased) in isolation, completely miss the true positive rate. On the other hand, the unbiased survey (k = 1) tracks the true positive rate well. The proposed method (magenta) also tracks the true positive rate, but the 95% equal-tailed credible intervals more tightly encapsulate the true positive rate.
@@
The reduction in credible interval width is quite noticeable. To quantify this, we take the mean ratio of the unbiased survey’s credible interval width (in isolation) to that of the proposed method. Across all time-points this mean ratio is 2.09, meaning the method cuts the credible interval width in half. This reduction is present to a lesser degree in the now-casting case. The credible interval for the last time-point using the unbiased survey is only 1.07 times larger than that of the synthesis method. Moreover, the method’s now-casted credible interval does contain the true positive rate whereas the unbiased survey’s interval does not.
@@ 3 Simulation study
To assess the extent to which the synthesis method reduces the now-casted MSE and the effects of mis-specifying the generating mechanism of $\phi_{kt}$, a 3 x 3 simulation design is performed. We consider 3 generating mechanisms for $\phi_{kt}$, where $\phi_{kt}$ either is constant in time (7), linear in time (8), or follows a random walk (9), all as described in the methods section. To get a good representation of how the methods perform under many parameter conditions we generate the parameters from the prior distributions. That is, each simulated dataset arises from different underlying parameter values. One issue from this technique is that priors presented in the methods section are wider than what might be realistic. For example, if one or more of the parameters are extreme in value, then $Y_{kt}$ may be equal to 0 or $n_{kt}$ for many values of $t$. To avoid this, we narrow the priors to more plausible values:
@@
$\sigma^{2}\sim\textrm{Normal}(0,0.1)T(0,\infty),$
@@
$\pi^{2}\sim\textrm{Normal}(0,0.01).$
This translates to $\textrm{logit}^{-1}(\theta_{0})$ having roughly 90% of its prior density between 0.22 and 0.78. Meaning most of the time we believe the initial positive rate is near 50%. We also set narrow our belief about $\sigma^{2}$ which means $\textrm{logit}^{-1}(\theta_{t})-\textrm{logit}^{-1}(\theta_{t-1})$ is $\pm$0.15 or less, 90% of the time, assuming $\textrm{logit}^{-1}(\theta_{t-1})=0.5$. For the linear model, $\phi_{kt}-\phi_{k(t-1)}$ in between -0.33 and 0.33 90% of the time when $\phi_{k(t-1)}=1$ Lastly, for the random walk model $\phi_{kt}-\phi_{k(t-1)}$ is between -0.37 and 0.40, 90% of the time, where $\phi_{k(t-1)}=1$. We simulate the data with $K=3$ surveys and with a population of $N=10,000,000$. Let $k=1$ be the unbiased survey, and $k=2$ and $3$ be the non-probability surveys. Given that non-probability surveys are often larger we set $n_{1t}=100$, and $n_{2t}=n_{3t}=1000$ for all $t$. We repeat this experiment for 3 sets of time-points: ($T$ = 5, 10, and 15). The models are all fit in JAGS 4.3.x (Plummer et al. 2003), with roughly 20,000-25,000 burn-in and a sample size of 50,000-70,000 depending on the complexity of $\phi_{kt}$. The goal is to compare the MSE, i.e., the root averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median across the 2000 parameters generated. We add 95% confidence intervals for the Monte Carlo MSE estimates based on the central limit theorem and the MCSE (Monte Carlo Standard Error), see Morris et al. (2019).
3.1 Simulation Issues
As mentioned before, the non-central hyper-geometric distribution (3) introduces computational difficulties. Therefore the models are fit using the binomial approximation (4), but the data are still generated from (3), as we believe this is the true data generating mechanism. With $N=10,000,000$ and the largest $n$ being $n_{kt}=1000$, the binomial approximation is adequate. Random non-central hyper-geometric values are generated using the MCMCpack library (Martin et al. 2011), which has an implementation based on work by Liao & Rosen (2001). Moreover, for simulations with a large number of time-points, $\phi_{kt}$ can grow quickly for large values of $t$ (if $\phi_{kt}$ is linear, (8)) and lead to computational issues. To fix this, the time-points, when the models are fitted, are centered in the $T=10$ and $15$ simulations. Meaning: $\phi_{kt}=\exp(\gamma_{k0}+\gamma_{kt}t^{\prime})$, where $t^{\prime}=t-\frac{T}{2}$.
@@ 3.2 Simulation Results
Figure 2: Results of simulation study for t = 5, 10, and 15 time-points. Simulations run using the model fit in JAGS 4.3.x. The MSE is calculated as the estimated averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median. Error bars represent the 95% confidence intervals for the Monte Carlo estimated MSE i.e. $\hat{\textrm{MSE}}\pm 1.96\textrm{MCSE}$.
Figure 2 shows the results of the simulation study. As expected, when the model for $\phi_{kt}$ is correctly specified, the MSE is the lowest. Moreover, incorporating the additional 2 biased surveys generally leads to a lower MSE, showing the information from the biased surveys is useful. However, if the model for $\phi_{kt}$ is drastically mis-specified, the synthesis model performs worse than using the unbiased survey in isolation. Specifically, fitting a constant $\phi_{kt}$, when $\phi_{kt}$ actually changes significantly, leads to a large increase in MSE. These simulations also suggest that the added complexity of fitting a random walk $\phi_{kt}$ does not lead to a large increase in estimator variance. Surprisingly, for scenarios with even a few time-points, a random walk model for $\phi_{kt}$ appears to be a relatively safe assumption. By increasing the number of time-points, additional complexity can be used to better adjust for the bias without gaining excessive estimator variance. Moreover, the confidence intervals of for Monte Carlo estimated MSE are sufficiently separated to have confidence in the results.
@@ 4 Large survey vaccine data
Monitoring the COVID-19 vaccination rates over time was an essential part of assessing the public health response to the pandemic. As discussed in the introduction, Bradley et al. (2021) explains how two survey were found to drastically overestimate the true vaccination rate. Out of the three surveys analyzed, only the Axios-Ipsos was found to track the CDC’s benchmark well (acknowledging imprecision in the benchmark). The Axios-Ipsos poll primarily focused on quality probabilistic sampling and maintained a high response rate (approx 50%). As a result, its confidence intervals were found to contain the CDC’s historically updated benchmark 10/11 times (up to June 2021), despite its small sample size. Thus, we propose using the method described in this paper to further improve the estimates of the Axios-Ipsos poll. We graciously use the extended data provided by Bradley et al. (2021) to assess the method on real-world data. We aim to compare the performance of the synthesis method to the Axios-Ipsos Poll and the CDC’s historically updated benchmark. We follow Bradley et al. (2021) in adding ±$\pm 5\%$ error to the CDC’s benchmark estimates. Let $\phi_{\textrm{AI}}=1$ for all $t$, and consider the three models for $\phi_{\textrm{DF}}$ and $\phi_{\textrm{HP}}$, where AI, DF, and HP refer to Axios-Ipsos, Delphi-Facebook, and Household-Pulse respectively. We consider two adjustments to the default prior settings suggested in Section 2. Firstly, $\theta_{0}\sim\textrm{Normal}(-2,1)$. Which means: roughly 90% of the mass of $\textrm{logit}^{-1}(\theta_{0})$ lies between 0.036 and 0.327. This reflects the prior belief in a low vaccination rate at time zero before seeing any data. Secondly, $\theta_{t}|\theta_{t-1}\sim\textrm{Normal}(\theta_{t-1},\sigma^{2})\textrm{T}(\theta_{t-1},\infty)$ which represents the knowledge that the vaccination rate can only increase.
@@ 4.2 Inference
In Figure 4 we see that the Axios-Ipsos survey alone has a larger 95% credible intervals for all time-points. The average and median Axios-Ipsos credible interval widths are 1.54 and 1.42 times larger respectively (where the Axios-Ipsos survey is not missing). This demonstrates the clear advantage of synthesising the information from multiple surveys, even if we acknowledge that the bias may be changing over time. The unbiasedness property of the Axios-Ipsos survey also appears to be perserved. The synthesis method’s 95% credible intervals are within the CDC’s benchmark (including the assumed CDC’s 5% margin of error) $43/46$ (93.5%) of the time (this excludes the last two time-points where no CDC data is present).
@@ 4.3 Now-cast performance
As well as making inference for each time-point after collecting the data on all time-points we also consider the now-casting performance. That is we only consider the information up to and including each time-point. In this section, we compare the now-cast point estimate and 95% credible interval for each of the 48 time-points. Each of the Axios-Ipsos, Houshold-Pulse, and Delphi-Facebook surveys are fit individually using the specified models in the methods section with $\phi_{\cdot t}=1$, for all $t$, and our method is fit using information from the three surveys up to and including each point with a random walk assumption for $\phi_{DF}$ and $\phi_{HP}$.
Figure 6: Plot of now cast performance of synthesis method. Estimates are posterior medians and equal-tailed 95% credible intervals. Figure extends data from Bradley et al. (2021)
In Figure 6 we can see the point estimates are similar the inference case (Figure 3), but with a considerable amount of added uncertainty. Again, the uncertainty of the synthesis method is much lower than that of the Axios-Ipsos survey used in isolation. Precisely, over the time-points where the Axios-Ipsos poll is not missing, the mean and median ratio of the Axios-Ipsos credible interval widths to that of the proposed method are both 1.24 times larger. This decrease in uncertainty does not appear to reduce the empirical credible interval coverage. The synthesis method’s 95% credible intervals intersect with the CDC’s benchmark range $44/46$ (95.7%) of the time. Moreover, there are a few time-points at which the method yields extremely large credible intervals, such as time-points 2 and 6. These are a result of the missing values in the unbiased data. Especially at the beginning of the time-points when few Axios-Ipsos surveys are seen, a missing value leads to a relatively wide estimation of $P(Y_{kt}|.)$, which ultimately increases the uncertainty about $\textrm{logit}^{-1}(\theta_{t})$.
@@
Figure 7 shows the posterior median and 95% credible intervals for $\sigma^{2}$ and $\pi^{2}$ across all 48 time-points. Note that for time-point 1, the estimate of $\sigma^{2}$ is essentially equal to that given by the prior. Around mid September there is a slight drop in the estimate for $\sigma^{2}$. This is because as the number of individuals who have been vaccinated starts to level off for both the Delphi-Facebook and Household-Pulse surveys.

### v3

@@ 1 Introduction
If we draw our attention to the similarities between causal inference in observational studies and non-representative sampling, as suggested by Mercer et al. (2017), we can better understand the main assumptions required to obtain unbiased estimates. The key assumption for post-survey adjustments is that of “conditional ignorability” (Mercer et al. 2017, Schuessler & Selb 2023). Much like the assumption of “no unmeasured confounders” that is required to obtain unbiased causal estimates in an observational study, the conditional ignorability assumption cannot be guaranteed. In other words, there is no sure way of choosing a sufficient set of adjustment covariates (based on available data alone), and ultimately there is no test or way to validate when the assumption of conditional ignorability is met.
@@
In Section 2, we outline the proposed method, state all assumptions, and provide an example where including biased surveys will reduce the uncertainty of estimates. In Section 3, we conduct a simulation study to show that using conservative assumptions, over-parametrizing the model is not overly detrimental, as long as the complexity of the model for the bias is proportional to the number of time-points. In Section 4, inspired by the dataset presented by Bradley et al. (2021), we show that the proposed method performs well in estimating the vaccination rate by combining two “biased” surveys with a presumably “unbiased” survey. The results show a noticeable reduction in uncertainty (compared to only using the unbiased survey). We conclude in Section 5 by discussing potential uses and limitations for using large amounts of unrepresentative survey data to improve the precision of survey estimates.
2 Method
@@
The hyper-geometric distribution describes the probability of drawing $Y_{kt}$ positive individuals amongst a sample of $n_{kt}$ who responded to the survey (without any individuals responding more than once), from a finite population of size $N$ that contains exactly $P_{t}$ positive individuals. The non-central hyper-geometric is a generalization of the hyper-geometric distribution whereby, in this case, the response may be biased so that either positive or negative individuals are more likely to respond. When $\phi_{kt}>1$, positive individuals are more likely to respond than negative individuals; when $\phi_{k}<1$, positive individuals are less likely to respond than negative individuals. When $\phi_{k}=1$, we have that the probability of responding is equal for both positive and negative individuals, and the non-central hyper-geometric distribution reduces to the standard hyper-geometric distribution. In this parameterization, the $\phi_{kt}$ parameter can be interpreted as an odds ratio: the odds of a positive individual being sampled vs. the odds of a non-positive individual being sampled. For a more detailed breakdown of the non-central hyper-geometric distribution, its use in population size estimation using surveys, and a discussion of Bayesian estimation of its parameters see Ballerini & Liseo (2022). As discussed in the introduction, there may be several factors that lead to positive individuals being sampled, many of them may be completely unknown to the practitioner which ultimately makes survey weighting difficult. In this model, we assume there is a multitude of factors that contribute to selection bias and that we need not directly measure them. $\phi_{kt}$ is the resulting odds ratio of the unknown function of all of these factors.
@@
To illustrate the proposed method, consider the goal of performing inference for $T=10$ time-points and $K=3$ surveys, with a population of $N=10,000$. Survey 1 is unbiased, whereas survey 2 and 3 are compromised by selection bias. Specifically, for these two surveys we generate data under the assumption of a linear $\phi_{kt}$ (8). All the parameters and data are generated jointly using (1),(2) (3), and (5), (6), (8) and the priors used in analysing the dataset are as such. The data are illustrated in Table 1.
@@
Analysis of each survey in isolation assumes $\phi_{kt}=1$ for all $t$, but retain all the same priors (1), (2), (3), and (5), (6). In Figure 1, we see that using the biased surveys ($k=2$ 3, assuming they are unbiased) in isolation, completely miss the true positive rate. On the other hand, the unbiased survey ($k=1$) tracks the true positive rate well. The proposed method (magenta) also tracks the true positive rate, but the 95% equal-tailed credible intervals more tightly encapsulate the true positive rate.
@@
The reduction in credible interval width is quite noticeable. To quantify this, we take the mean ratio of the unbiased survey’s credible interval width (in isolation) to that of the proposed method. Across all time-points this mean ratio is 2.18, meaning the method cuts the credible interval width in half. This reduction is present to a lesser degree in the now-casting case. The credible interval for the last time-point using the unbiased survey is only 1.12 times larger than that of the synthesis method. Moreover, the method’s now-casted credible interval does contain the true positive rate whereas the unbiased survey’s interval does not.
@@ 3 Simulation study
To assess the extent to which the synthesis method reduces the now-casted MSE and the effects of over-parametrizing the generating mechanism of $\phi_{kt}$, a 3 x 3 simulation design is performed. We consider 3 generating mechanisms for $\phi_{kt}$, where $\phi_{kt}$ either is constant in time (7), linear in time (8), or follows a random walk (9), all as described in the methods section. To get a good representation of how the methods perform under many parameter conditions we generate the parameters from the prior distributions. That is, each simulated dataset arises from different underlying parameter values. One issue from this technique is that priors presented in the methods section are wider than what might be realistic. For example, if one or more of the parameters are extreme in value, then $Y_{kt}$ may be equal to 0 or $n_{kt}$ for many values of $t$. To avoid this, we narrow the priors to more plausible values:
@@
$\sigma^{2}\sim\textrm{Normal}(0,0.1)\textrm{T}(0,\infty),$
@@
$\pi^{2}\sim\textrm{Normal}(0,0.01)\textrm{T}(0,\infty).$
This translates to $\textrm{logit}^{-1}(\theta_{0})$ having roughly 90% of its prior density between 0.22 and 0.78. Meaning most of the time we believe the initial positive rate is near 50%. We also set narrow our belief about $\sigma^{2}$ which means $\textrm{logit}^{-1}(\theta_{t})-\textrm{logit}^{-1}(\theta_{t-1})$ is $\pm$0.15 or less, 90% of the time, assuming $\textrm{logit}^{-1}(\theta_{t-1})=0.5$. For the linear model, $\phi_{kt}-\phi_{k(t-1)}$ in between -0.33 and 0.33 90% of the time when $\phi_{k(t-1)}=1$ Lastly, for the random walk model $\phi_{kt}-\phi_{k(t-1)}$ is between -0.37 and 0.40, 90% of the time, where $\phi_{k(t-1)}=1$. We simulate the data with $K=3$ surveys and with a population of $N=10,000,000$. Let $k=1$ be the unbiased survey, and $k=2$ and $3$ be the non-probability surveys. Given that non-probability surveys are often larger we set $n_{1t}=100$, and $n_{2t}=n_{3t}=1000$ for all $t$. We repeat this experiment for 3 sets of time-points: ($T$ = 5, 10, and 15). The models are all fit with 10 parallel chains using the dclone package (Sólymos 2010) which links to JAGS 4.3.x (Plummer et al. 2003). Each chain has roughly 20,000-25,000 burn-in iterations and 50,000-70,000 draws depending on the complexity of $\phi_{kt}$, all with a thinning interval of 5. The goal is to compare the MSE, i.e., the averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median across the 2000 parameters generated. We add 95% confidence intervals for the Monte Carlo MSE estimates based on the central limit theorem and the MCSE (Monte Carlo Standard Error), see Morris et al. (2019).
3.1 Simulation considerations
As mentioned before, the non-central hyper-geometric distribution (3) introduces computational difficulties. Therefore the models are fit using the binomial approximation (4), but the data are still generated from (3), as we believe this is the true data generating mechanism. With $N=10,000,000$ and the largest $n$ being $n_{kt}=1000$, the binomial approximation is adequate. Random non-central hyper-geometric values are generated using the MCMCpack library (Martin et al. 2011), which has an implementation based on work by Liao & Rosen (2001). Random values from the truncated normal are generated via the truncnorm library (Mersmann et al. 2018), and the rest of values are generated using R (R Core Team 2023). Moreover, for simulations with a large number of time-points, $\phi_{kt}$ can grow quickly for large values of $t$ (if $\phi_{kt}$ is linear, (8)) and lead to computational issues. To fix this, the time-points, when the models are fitted, are centered in the $T=10$ and $15$ simulations. Meaning: $\phi_{kt}=\exp(\gamma_{k0}+\gamma_{kt}t^{\prime})$, where $t^{\prime}=t-\frac{T}{2}$.
@@ 3.2 Simulation Results
Figure 2: Results of simulation study for t = 5, 10, and 15 time-points. Simulations run using the model fit in JAGS 4.3.x. The MSE is calculated as the estimated averaged squared difference between $\textrm{logit}^{-1}(\theta_{T})$ and the posterior median. Error bars represent the 95% confidence intervals for the Monte Carlo estimated MSE i.e. $\widehat{\textrm{MSE}}\pm 1.96\textrm{MCSE}$.
Figure 2 shows the results of the simulation study. As expected, when the model for $\phi_{kt}$ is correctly specified, the MSE is the lowest. Moreover, incorporating the additional 2 biased surveys generally leads to a lower MSE, showing the information from the biased surveys is useful. However, if the model for $\phi_{kt}$ is drastically over-parametrized, the synthesis model performs worse than using the unbiased survey in isolation. Specifically, fitting a constant $\phi_{kt}$, when $\phi_{kt}$ actually changes significantly, leads to a large increase in MSE. These simulations also suggest that the added complexity of fitting a random walk $\phi_{kt}$ does not lead to a large increase in estimator variance. Surprisingly, for scenarios with even a few time-points, a random walk model for $\phi_{kt}$ appears to be a relatively safe assumption. By increasing the number of time-points, additional complexity can be used to better adjust for the bias without gaining excessive estimator variance. Moreover, the confidence intervals of for Monte Carlo estimated MSE are sufficiently separated to have confidence in the results.
@@ 4 Large survey vaccine data
Monitoring the COVID-19 vaccination rates over time was an essential part of assessing the public health response to the pandemic. As discussed in the introduction, Bradley et al. (2021) explains how two survey were found to drastically overestimate the true vaccination rate. Out of the three surveys analyzed, only the Axios-Ipsos was found to track the CDC’s historically updated benchmark well (acknowledging imprecision in the benchmark). The Axios-Ipsos poll primarily focused on quality probabilistic sampling and maintained a high response rate (approx 50%). As a result, its confidence intervals were found to contain the CDC’s historically updated benchmark 10/11 times (up to June 2021), despite its small sample size. Thus, we propose using the method described in this paper to further improve the estimates of the Axios-Ipsos poll. We graciously use the extended data provided by Bradley et al. (2021) to assess the method on real-world data. We aim to compare the performance of the synthesis method to the Axios-Ipsos Poll and the CDC’s historically updated benchmark. We follow Bradley et al. (2021) in adding ±$\pm 5\%$ error to the CDC’s benchmark estimates. Let $\phi_{\textrm{AI}}=1$ for all $t$, and consider the three models for $\phi_{\textrm{DF}}$ and $\phi_{\textrm{HP}}$, where AI, DF, and HP refer to Axios-Ipsos, Delphi-Facebook, and Household-Pulse respectively. We consider two adjustments to the default prior settings suggested in Section 2. Firstly, $\theta_{0}\sim\textrm{Normal}(-2,1)$. Which means: roughly 90% of the mass of $\textrm{logit}^{-1}(\theta_{0})$ lies between 0.036 and 0.327. This reflects the prior belief in a low vaccination rate at time zero before seeing any data. Secondly, $\theta_{t}|\theta_{t-1}\sim\textrm{Normal}(\theta_{t-1},\sigma^{2})\textrm{T}(\theta_{t-1},\infty)$ which represents the knowledge that the vaccination rate can only increase.
@@ 4.2 Inference
In Figure 4 we see that the Axios-Ipsos survey alone has a larger 95% credible intervals for all time-points. The average and median Axios-Ipsos credible interval widths are 1.54 and 1.42 times larger respectively (where the Axios-Ipsos survey is not missing). This demonstrates the clear advantage of synthesising the information from multiple surveys, even if we acknowledge that the bias may be changing over time. The unbiasedness property of the Axios-Ipsos survey also appears to be preserved. The synthesis method’s 95% credible intervals are within the CDC’s historically updated benchmark (including the assumed CDC’s 5% margin of error) $43/46$ (93.5%) of the time (this excludes the last two time-points where no CDC data is present).
@@ 4.3 Now-cast performance
As well as making inference for each time-point after collecting the data on all time-points we also consider the now-casting performance. That is we only consider the information up to and including each time-point. In this section, we compare the now-cast point estimate and 95% credible interval for each of the 48 time-points. Each of the Axios-Ipsos, Household-Pulse, and Delphi-Facebook surveys are fit individually using the specified models in the methods section with $\phi_{\cdot t}=1$, for all $t$, and our method is fit using information from the three surveys up to and including each point with a random walk assumption for $\phi_{DF}$ and $\phi_{HP}$.
Figure 6: Plot of now cast performance of synthesis method. Estimates are posterior medians and equal-tailed 95% credible intervals. Figure extends data from Bradley et al. (2021).
In Figure 6 we can see the point estimates are similar the inference case (Figure 3), but with a considerable amount of added uncertainty. Again, the uncertainty of the synthesis method is much lower than that of the Axios-Ipsos survey used in isolation. Precisely, over the time-points where the Axios-Ipsos poll is not missing, the mean and median ratio of the Axios-Ipsos credible interval widths to that of the proposed method are both 1.24 times larger. This decrease in uncertainty does not appear to reduce the empirical credible interval coverage. The synthesis method’s 95% credible intervals intersect with the CDC’s historically updated benchmark range $44/46$ (95.7%) of the time. Moreover, there are a few time-points at which the method yields extremely large credible intervals, such as time-points 2 and 6. These are a result of the missing values in the unbiased data. Especially at the beginning of the time-points when few Axios-Ipsos surveys are seen, a missing value leads to a relatively wide estimation of $P(Y_{kt}|.)$, which ultimately increases the uncertainty about $\textrm{logit}^{-1}(\theta_{t})$.
@@
Figure 7 shows the posterior median and 95% credible intervals for $\sigma^{2}$ and $\pi^{2}$ across all 48 time-points. Note that for time-point 1, the estimate of $\sigma^{2}$ is essentially equal to that given by the prior. Around mid September there is a slight drop in the estimate for $\sigma^{2}$. This is because as the number of individuals who have been vaccinated starts to level off for both the Delphi-Facebook and Household-Pulse surveys. The uncertainty around $\pi^{2}$ also reduces quite quickly. After only about 7-8 time-points most of belief of the variance in the movement in $\phi_{kt}$ has been narrowed.